in a triangle, one angle is twice as large as another angle and the third angle is 40 degree larger than the smallest angle . find all three angles.
Answer by Cromlix(4381)

First angle = x
Second angle = 2x
Third angle = x + 40
Total angles in a triangle = 180 degrees
x + 2x + x + 40 = 180
Collect like terms
4x + 40 = 180
4x = 180 - 40
4x = 140
x = 35 degrees
First angle = 35 degrees
Second angle = 70 degrees
Third angle = 75 degrees
Hope this helps.
:-)